# Sparse Repellency for Shielded Generation in Text-to-Image Diffusion Models

### v1

@@ 4.2 Benchmark
We first examine whether adding SPELL post-hoc increases the diversity of trained diffusion models. To this end, we quantitatively compare each diffusion model to the same model run with the same random generation seeds but with SPELL. Table 1 shows that SPELL consistently increases the diversity in terms of both recall and Vendi score across all models, in both text-to-image and class-to-image diffusion. The third diversity metric, coverage, is also increased in all models except SD3. The difference between coverage and recall is that coverage uses a more tight neigborhood radius to determine whether an image of the original dataset is covered by the generated ones. However, SD3 is the only method that was not trained on ImageNet-1k/CC12M. We find that SPELL correctly helps SD3 generate images that are generally more diverse, as evidenced by the 26% and 37% increases in the reference-dataset-free Vendi score, but in other attributes than in the reference datasets, explaining the decrease in coverage. Precision and density decrease slightly in 5 out of the six models, but stay within -3% to +2% of the original values, the exception again being SD3. This tradeoff between diversity and precision is common in the literature (Kynkäänniemi et al., 2024; Sadat et al., 2024; Corso et al., 2024), and we show in the next section that our method provides more favorable Pareto fronts than alternative recent methods. This tradeoff improves the overall $\text{FD}_{\text{DINOv2}}$ score considerably in Latent Diffusion and MDTv2, while staying within 3% of the original value on Simple Diffusion and EDMv2, and increasing on SD3. Overall, we find that SPELL increases the diversity considerably across all models, with only minor tradeoffs in precision.
@@ 4.3 Comparison to other diversity-inducing methods
We now take a closer look at SPELL’s hyperparameter, the repellency radius $r$. Figure 3 shows that we can use repellency in two ways. A low radius increases the diversity (Vendi score and recall) without compromising on precision. This small radius intuitively prevents generating the nearly-same image twice. The resulting output set also better reflects the true image distribution, yielding an enhanced $\text{FD}_{\text{DINOv2}}$. The second option is to further increase the repellency radius, gaining diversity at the cost of precision.
@@
Figure 4 shows that SPELL achieves a more favorable trade-off curve than each of these methods, in three different diversity vs quality Pareto fronts, namely recall vs precision, coverage vs density, and Vendi score vs CLIP Score. One reason for this is SPELL’s sparsity. SPELL only intervenes if images are going to be too similar to one another. Due to it’s ReLU weighting, it remains exactly zero if an image’s trajectory is already headed to a diverse output. This leads to increased diversity while leaving high-precision images unchanged. We study this sparsity further in the following section.
@@ 4.4 Sparsity analysis
Figure 5: (a) The gradient that SPELL adds is only a fraction of the magnitude of the diffusion’s score, thus adjusting it without drowning it out. (b) Repellency happens primarily in early backwards steps and then remains zero, thus making it sparse. Overcompensation allows finishing the repellency earlier, whereas runs with high repellency radius repel longer. Latent Diffusion trajectories with repellency from both previous generations and intra-batch.
@@ 4.5 Qualitative Example of SPELL Interventions
Figure 6: Images for the prompt “a dog plays with a ball” generated one by one, $B=1$, with SD3 (top) and SD3 + SPELL (bottom), using the same random seeds. SPELL intervenes on 10 of the 16 images, marked with orange borders, because they are too similar to previously generated images. Images that SPELL repelled from are listed below each image. Early images are adjusted less often and mostly by details because they are still novel enough. Later images repel from more previous images and more strongly to ensure they are different enough.
@@ 4.6 Image protection benchmark
Table 2: SPELL almost always generates images outside the shields of ImageNet-1k train set. The number of Voronoi cells searched during nearest neighbor lookup controls the tradeoff between protection rate and inference-time compute. The runtime is reported on a single A100-40GB GPU.
Model Searched cells Unprotected images $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Time per image (s) $\downarrow$
@@ Appendix B Repellence Guarantee
Appendix C Conservative field
@@ Appendix E Construction of the soft-label CC12M dataset
Appendix F Ablation: Repellency throughout training

### v2

@@ 4.2 Benchmark
We first examine whether adding SPELL post-hoc increases the diversity of trained diffusion models. To this end, we quantitatively compare each diffusion model to the same model run with the same random generation seeds but with SPELL. In particular, we use intra-batch repellency together with repellency from previously generated batches, to enable repellency across the up to 128 images per prompt/class. Table 1 shows that SPELL consistently increases the diversity, both in terms of recall and Vendi score, across all text-to-image and class-to-image diffusion models. This demonstrates that SPELL works independently from the model architecture and the space the models diffuse in (RGB space for Simple Diffusion, VAE space for all others). Coverage remains within -1% to +2% of its original value in all models except SD3. The difference between coverage and recall is that coverage uses a more tight neigborhood radius to determine whether an image of the original dataset is covered by the generated ones. In other words, coverage measures a form of dataset match, which can counter-intuitively be decreased by more diverse outputs if the diversity takes different forms or is higher than in the reference dataset. This stands out most for SD3, which was not trained on the reference datasets ImageNet-1k/CC12M. We find that SPELL correctly helps SD3 generate images that are generally more diverse, as evidenced by the 26% and 37% increases in the reference-dataset-free Vendi score, but in other attributes than in the reference datasets, explaining the decrease in coverage. Precision and density decrease slightly in five out of the six models, but stay within -3% to +2% of the original values, the exception again being SD3. This tradeoff between diversity and precision is common in the literature (Kynkäänniemi et al., 2024; Sadat et al., 2024; Corso et al., 2024), and we show in the next section that our method provides more favorable Pareto fronts than alternative recent methods. This tradeoff improves the overall $\text{FD}_{\text{DINOv2}}$ score considerably in Latent Diffusion and MDTv2, while staying within 3% of the original value on Simple Diffusion and EDMv2, and increasing on SD3. Overall, we find that SPELL increases the diversity considerably across all models, with only minor tradeoffs in precision.
@@ 4.3 Comparison to Other Diversity-inducing Methods
We now take a closer look at SPELL’s hyperparameter, the repellency radius $r$. Figure 3 shows that we can use repellency in two ways. A low radius ($r=15$) increases the diversity (Vendi score and recall) without compromising on precision. Intuitively, such a small radius only serves to prevent generating the nearly-same image twice. The resulting output set also better reflects the true image distribution, yielding an enhanced $\text{FD}_{\text{DINOv2}}$. The second option is to further increase the repellency radius, which as Figure 3 shows further enhances diversity at the cost of precision.
@@
Figure 4 shows that SPELL achieves a more favorable trade-off curve than each of these methods, in three different diversity vs quality Pareto fronts, namely recall vs precision, coverage vs density, and Vendi score vs CLIP Score. One reason for this is SPELL’s sparsity. SPELL only intervenes if images are going to be too similar to one another. Due to its ReLU weighting, it remains exactly zero if an image’s trajectory is already headed to a diverse output. This leads to increased diversity while leaving high-precision images unchanged. We study this sparsity further in the following section.
@@ 4.4 Sparsity Analysis
Figure 5: (a) The gradient that SPELL adds is only a fraction of the magnitude of the diffusion’s score, thus adjusting it without drowning it out. (b) Repellency happens primarily in early backwards steps ($t\in[0.6,1.0])$ and then often remains zero, thus making it sparse. Overcompensation allows finishing the repellency earlier, whereas runs with high repellency radius repel longer. Latent Diffusion trajectories with repellency from both previous generations and intra-batch.
@@ 4.5 Qualitative Example of SPELL Interventions
Figure 6: Images for the prompt “a dog plays with a ball” generated one-by-one ($B=1$) with SD3 (top) and SD3 + SPELL (bottom), using the same random seeds. SPELL intervenes on 10 of the 16 images, marked with orange borders, because they are too similar to previously generated images. Images that SPELL repelled from are listed below each image. Early images are adjusted less often and mostly in details because they are still novel enough. Later images repel from more previous images and more strongly to ensure they are different enough.
@@ 4.6 Image Protection Benchmark
Table 2: SPELL almost always generates images outside the shields of the ImageNet-1k train set. Searching for shields in more Voronoi cells during the nearest neighbor lookup improves the protection rate at a higher inference-time runtime. The runtime is reported on a single A100-40GB GPU.
Model Searched cells Generated images too close to ImageNet neighbors $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ Time per image (s) $\downarrow$
@@ Appendix B Repellence Guarantee
Appendix C Conservative Field Interpretation
@@ Appendix E Construction of the Soft-label CC12M Dataset
Appendix F Ablation: Repellency Strength Throughout the Generation